OpenStudy (anonymous):

I want to construct a nice Baye's theorem math problem and make a video. I need ideas. It should be super counter intuitive and a physical one. I just thought of one "Did she break up with you?"

OpenStudy (anonymous):

Suppose you are in a long term relation with your girl. Super awesome So the probability of she wanting to break up is just lets say 1% Now we all get annoyed when our partners don't receive your calls. lets assign some probabilities If your girl wants to break up with you, there is a 90% chance that she wouldn't wanna talk to you at all, and she won't receive your call However when that is not the case, when your girl is busy maybe doing something else, there is a 10% chance that she wouldn't receive your call Now you call her up, and she DOESN"T RECEIVE IT What is the probability that she wnts to break up with you? :D

OpenStudy (anonymous):

Is this a nice question or is it boring?! :-/ or any ideas?

OpenStudy (anonymous):

@Kainui what you think? :P

OpenStudy (anonymous):

@ganeshie8 @myininaya @iambatman @Compassionate @nincompoop @bibby

OpenStudy (kainui):

Hahaha I think this is an interesting idea, I like it a lot!

OpenStudy (anonymous):

Thank you @Kainui I worked out the probability to be 8.3% ... can u confirm that? :D

OpenStudy (kainui):

I'm not gonna lie, Bayes theorem is my weakness. I just can't ever seem to understand it because it always seems counter intuitive to me.

OpenStudy (anonymous):

Lol.. :D.. all Hail Bayes!!

OpenStudy (kainui):

Yeah I was hoping you could explain it to me actually lol.

OpenStudy (anonymous):

I can do the formula:P.. But I m still working on an intuitive approach.. :P

OpenStudy (mathmate):

Yes, I get 0.08333 as well !

OpenStudy (anonymous):

yes thank you :) :)

OpenStudy (anonymous):

:P

OpenStudy (mathmate):

You're welcome! I think you have an attractive story to introduce the subject!

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

P("yes for breakup" | "didn't receive the call") = P("yes for breakup" and "didnt recieve the call") / P("didnt receuve the call") = (0.01*0.9) / (0.01*0.9 + 0.99*0.1) =0.0833 for more accurate inference, i think you will need to repeat the experiment by calling her multiple times, each time updating the probabilities =P

ganeshie8 (ganeshie8):

this is a good example to introduce bayee's theorem, good luck !

OpenStudy (anonymous):

it reminds me of odd logistic regression

OpenStudy (anonymous):

@ganeshie8 I am not actually doing any series on theorem, I m just making random cool videos , thought I ll do one on this, just to get people curious about this awesome theorem, making them think as to "The froockk.. how did that happen" :D

OpenStudy (anonymous):

bhaha theory of awesomeness